OpenStudy (anonymous):

Find the domain and range of Y=x^2-9 and graph the function.

OpenStudy (jdoe0001):

well, what values do you think you can give "x"?

OpenStudy (debbieg):

It's a quadratic, and quadratics have: domain = all reals range = [min, infinity) if upward opening, or range = (-infinity, max] if downward opening

OpenStudy (anonymous):

Domain: All Real Numbers Range: [-9, Inf)

OpenStudy (anonymous):

Sorry guys I was in a car ride. to jdoe0001. I know the domain must be all real numbers because there are no restrictions on the x value. I do not know how to find the range though

OpenStudy (jdoe0001):

well, if you notice any negative values thrown in to x^2 will go like this \(\bf x^2\\ (-1)^2 \implies -1 \times -1 =1\\ (-10)^2 \implies -10 \times -10 =100\\ (-100)^2 \implies -100 \times -100 =10000\) so you see, no matter what negative values you'd give to x^2, it will always throw back out a positive number what's the lowest "x" can get? well, x= 0 => (0)^2 = 0 at that point \(\bf y=x^2-9 \implies y=(0)^2-9 \implies y = 0-9\implies y = -9\) so, that's the lowest "y" will get, so it's range as will be no lower than -9 and up to infinity, or as boricua86 said, \(\bf [- 9, \infty)\)

OpenStudy (anonymous):

OHH that was a great way to explain it thank you! I should be able to figure the rest of these types of problems out now. The graph is just a parabola shifted downward 9 right?

OpenStudy (jdoe0001):



OpenStudy (anonymous):

thanks!

OpenStudy (jdoe0001):

yw